In a case in Riverhead, New York, nine different crime victims listened to voice recordings of five different men. All nine victims identified the same voice as that of the criminal. If the voice identifications were made by random guesses, find the probability that all nine victims would select the same person. Does this constitute reasonable doubt?
1
Expert's answer
2012-09-13T11:35:53-0400
Let the first victim identified the voice of person A.
The probability of
selecting the same person is 1/5, repeated by eight victims:
P = (1/5)^8 = 1
/ (5^8) = 1 / 390625.
So, there is no reasonable doubt.